# Quantifying Phase Unbalance and Coordination Impacts on Distribution Network Flexibility

### v1

@@ III Case Studies
In this section, the proposed framework is demonstrated using two case studies: (i) a simple 5-bus system with DER flexibility available at a single location; (ii) a real 221-bus distribution network in the UK with 12 flexible units. Data for these cases is available at [38] and is also included in the code repository [28]. For each system, the limits of flexibility services are characterised via the concept of aggregated P-Q flexibility areas. Then, the impacts of voltage unbalance and phase coordination are quantified and the main findings are discussed. All simulations have been performed with JuMP 1.11.0 for Julia 1.6.1 language and Ipopt 3.14.4 solver.
@@ III-A Case Study: Illustrative 5-bus System
To illustrate the principles of flexibility services estimation considering phase unbalance, a simple network with 5 buses is constructed, visualised in Fig. 2. The network has four loads and one flexibility provider. Each load has connections to three phases, which enables to simulate balanced network operation or to create phase unbalance due to uneven allocation of single-phase loads. [^2] In this work, it is assumed that a three-phase balanced voltage source is connected to the primary substation, representing power supply from a medium voltage or high voltage network. That is, the voltage of the source is set to 1.0 p.u. for each phase. The inclusion of unbalanced power supply sources is left for future research. The flexible unit also has connections to all phases and therefore can be used to simulate balanced or unbalanced flexible power output.
Specifically, the following network conditions are considered in simulations for this case study:
@@
First, the impact of load unbalance on flexibility services is analysed. The flexible unit is assumed to be a three-phase device producing the same flexible power output per phase. That is, by design, this unit can provide only balanced flexibility services. The question is, does the load unbalance, inherent in LV distribution networks, affect such services? To address this question, the aggregated P-Q flexibility area of the network was calculated for cases with balanced and unbalanced loads, as shown in Fig. 3. Each boundary was approximated by 80 points, i.e., the flexibility estimation model (1a)-(1p) was iteratively solved 80 times. In the case of a balanced network, the flexibility service is less constrained and has the largest P-Q area (displayed in grey colour). [^3] Note that for a perfectly balanced network, there is no need to display flexibility areas per each phase as these areas coincide. In the case of an unbalanced network, the operation of the flexible unit is constrained differently in each phase. Therefore, the limits of the flexibility service are reduced, and the P-Q areas become different in the phases (as highlighted in blue, orange, and green colours). The quantification of the flexibility reduction due to load unbalance is given in Table I, where the areas of aggregated flexibility are calculated in kVAr2. The results demonstrate that the areas are significantly reduced due to the introduced load unbalance (36-38% reduction for the 5-bus system). It follows that balanced flexibility services provided by three-phase devices can be constrained by load unbalance present in distribution networks.
@@
Second, the impact of voltage unbalance limits on flexibility services is analysed. An unbalanced network with three single-phase flexible units is considered. That is, both loads cause phase unbalance and the units can provide unbalanced flexibility services. Then, the VUF constraints (1l)-(1o) are imposed to control voltage unbalance caused by the uneven allocation of single-phase loads and unbalanced output of single-phase flexible units. The aggregated P-Q flexibility areas for VUF in the range 0.1-1.5% are shown in Fig. 4.444Note that VUF limits below 1% are unrealistically low – no grid codes regulate such low voltage unbalances. In this work, low VUF limits are included in the analysis for the sake of illustrating the principles of flexibility provision under phase unbalance. [^5] In the 5-bus system, VUF constraints (1l)-(1o) are imposed on each bus. For simplicity, only phase $\phi=\{b\}$ is displayed in the figure – flexibility services in other phases have similar dependencies on VUF constraints.
@@
However, for current real distribution systems, the perfect coordination assumption may be unreasonable since 1) single-phase flexible units can be connected at different locations and belong to different owners and 2) required automation and communication technologies are not widely adopted. To analyse the lack of unit coordination, phase coordination constraint (1p) is imposed. Under this constraint, flexibility service in phase $\phi$ can be provided only by the single-phase unit connected to this phase. Other units cannot manage the voltage unbalance or assist in maximising network flexibility. The impact of phase coordination on P-Q flexibility areas is shown in Fig. 5, where VUF constraints (1o) in the range 0.1-1.0% are iteratively imposed. The findings of this simulation are twofold: (i) without coordination between units connected to different phases, the limits of aggregated flexibility reduce significantly, for example, 60.76% flexibility area reduction for phase $\phi=\{b\}$; (ii) voltage unbalance limits have a major impact on flexibility services provided by units with a lack of coordination. It follows that a lack of phase coordination combined with voltage unbalance limits creates the worst conditions for providing flexibility services: single-phase units cannot effectively manage voltage unbalance across the network, which results in the infeasibility of services provision.
@@ III-B Case Study: 221-bus LV Distribution Network in the UK
To investigate the impact of phase unbalance on the network’s flexibility services, 12 single-phase flexible units (four units per phase) were placed in the network as shown in Fig. 6. Each unit is assumed to have a P-Q capability of ±5 kW and ±5 kVAr. That is, units can independently regulate their active and reactive power output in the range of [-5.0,5.0] kW and [-5.0,5.0] kVAr. Then, the aggregated P-Q flexibility areas were calculated per each phase by iteratively solving the flexibility estimation model (1a)-(1p). The impact of phase unbalance and coordination on P-Q flexibility areas is shown in Fig. 7, where VUF constraints (1o) in the range 0.5-1.0% are imposed. [^6] To avoid scalability issues, in the 221-bus system, VUF constraints (1l)-(1o) are imposed only on the 7 buses with the highest voltage unbalances.
@@
Note that in this simulation setting, all flexible units have been assigned the same P-Q capabilities and distributed evenly across phases. Yet, their operation is limited considerably due to phase coordination issues and voltage unbalance. It follows that the provision of flexibility services by DER can be challenging even in LV networks with low load unbalance as flexible units can exacerbate existing voltage unbalances by producing or consuming power. To effectively manage such unbalances and voltage constraints, units connected to different phases must be perfectly coordinated. The simulations performed for the realistic 221-bus network demonstrate that lack of coordination between units significantly reduces the potential of DER flexibility services and can magnify voltage unbalance problems. Other aspects and implications of these findings are discussed in the next section.
@@ IV Discussion
Multiple aspects of flexibility provision in unbalanced LV distribution networks have to be further investigated. One of the key factors affecting the aggregated P-Q flexibility is the relationship between the placement of flexible units and locations of voltage unbalances. For instance, in the 221-bus case study displayed in Fig. 6, the highest load unbalance is happening at the feeder with the flexible unit F5{a} connected. Six more flexible units are placed in this part of the network. Therefore, the design of this experiment inherently creates constraints on the operation of these flexible units due to the nearby voltage unbalances. That is, units must be coordinated when consuming or producing power to manage voltage constraints and keep the network’s operation feasible. For example, units connected to phase $\phi=\{a\}$ (the phase with the highest load) must not increase their power consumption while units in other phases consume power. Otherwise, voltage unbalance in some parts of the network may increase to unacceptable levels or even the voltage limits (1b) for phase $\{a\}$ may be violated. Note that this would not happen if flexible units are located far from the load unbalance, say, at the primary substation. Therefore, the question arises: how does the initial network unbalance impact the network’s P-Q flexibility and how does it restrict flexible resources at different locations? This question has implications for DER planning models. Depending on DSO metrics and objectives, some DER should be placed closer to the primary substation, while other flexible resources should be located closer to existing voltage unbalances to manage voltage constraints there. A similar problem was considered in [23], where it was demonstrated that suboptimal selection of the DER connection phase can lead to voltage unbalance and limit DER capacity.
@@ V Conclusion
This paper investigates the provision of flexibility services in unbalanced LV distribution networks and proposes a framework for quantifying the impacts of phase unbalance and imperfect coordination between distributed flexibility providers. At the core of the framework lies an accurate nonlinear three-phase AC OPF model designed for simulating the operation of flexible units connected to different phases. To characterise feasible flexibility services available in networks, the concept of aggregated P-Q flexibility areas is applied. The combination of these models and concepts allows to directly translate voltage unbalance and phase coordination constraints as reductions of the aggregated P-Q flexibility. The simulations, performed for an illustrative 5-bus distribution system and a real 221-bus LV network in the UK, demonstrate that a significant share (over 30%) of total aggregated DER flexibility potential may be unavailable due to voltage unbalances and lack of coordination between DER connected to different phases.

### v2

@@ III Case Studies
In this section, the proposed framework is demonstrated using two case studies: (i) a simple 5-bus system with DER flexibility available at a single location; (ii) a real 221-bus distribution network in the UK with 12 flexible units. Data for these cases is available at [39, 29]. For each system, the limits of flexibility services are characterised via the concept of aggregated P-Q flexibility areas. Then, the impacts of voltage unbalance and phase coordination are quantified. Finally, a stochastic analysis is performed to evaluate how frequently voltage unbalance violations can occur due to uncoordinated flexible resources. All simulations have been performed with JuMP 1.11.0 for Julia 1.6.1 language and Ipopt 3.14.4 solver.
@@ III-A Case Study: Illustrative 5-bus System
To illustrate the principles of flexibility services estimation considering phase unbalance, a simple network with 5 buses is constructed (see Fig. 3). The network has four loads and one flexibility provider. Each load has connections to three phases, which enables to simulate balanced network operation or to create phase unbalance due to uneven allocation of single-phase loads. The flexible unit also has connections to all phases and therefore can be used to simulate balanced or unbalanced flexible power output.
The following network conditions are considered in simulations for this case study:
@@
First, the impact of load unbalance on flexibility services is analysed. The flexible unit is assumed to be a three-phase device producing the same flexible power output per phase. That is, by design, this unit can provide only balanced flexibility services. The question is, does the load unbalance, inherent in LV distribution networks, affect such services? To address this question, the aggregated P-Q flexibility area of the network was calculated for cases with balanced and unbalanced loads, as shown in Fig. 4. Each boundary was approximated by 80 points, i.e., the flexibility estimation model (1a)-(1p) was iteratively solved 80 times. In the case of a balanced network, the flexibility service is less constrained and has the largest P-Q area (displayed in grey colour). [^2] Note that for a perfectly balanced network, there is no need to display flexibility areas per each phase as these areas coincide. In the case of an unbalanced network, the operation of the flexible unit is constrained differently in each phase. Therefore, the limits of the flexibility service are reduced, and the P-Q areas become different across the three phases (as highlighted in blue, orange, and green colours). The quantification of the flexibility reduction due to load unbalance is given in Table I, where the areas of aggregated flexibility are calculated in kVAr2. The results demonstrate that the areas are significantly reduced due to the introduced load unbalance (36-38% reduction for the 5-bus system). It follows that balanced flexibility services provided by three-phase devices can be constrained by load unbalance present in distribution networks.
@@
Second, the impact of voltage unbalance limits on flexibility services is analysed. An unbalanced network with three single-phase flexible units is considered. That is, both loads cause phase unbalance and the units can provide unbalanced flexibility services. Then, the VUF constraints (1l)-(1o) are imposed to control voltage unbalance caused by the uneven allocation of single-phase loads and unbalanced output of single-phase flexible units. The aggregated P-Q flexibility areas for VUF in the range 0.1-1.5% are shown in Fig. 5.333Note that VUF limits below 1% are unrealistically low – no grid codes regulate such low voltage unbalances. In this work, low VUF limits are included in the analysis for the sake of illustrating the principles of flexibility provision under phase unbalance., [^4] In the 5-bus system, VUF constraints (1l)-(1o) are imposed on each bus. For simplicity, only phase $\phi=\{b\}$ is displayed in the figure – flexibility services in other phases have similar dependencies on VUF constraints.
@@
However, for existing distribution systems, the perfect coordination assumption may be unreasonable since 1) single-phase flexible units can be connected at different locations and belong to different owners, and 2) required automation and communication technologies are not widely adopted. To analyse the lack of unit coordination, phase coordination constraint (1p) is imposed. Under this constraint, flexibility service in phase $\phi$ can be provided only by the single-phase unit connected to this phase. Other units cannot manage the voltage unbalance or assist in maximising network flexibility. The impact of phase coordination on P-Q flexibility areas is shown in Fig. 6, where VUF constraints (1o) in the range 0.1-1.0% are iteratively imposed. The findings of this simulation are twofold: (i) without coordination between units connected to different phases, the limits of aggregated flexibility reduce significantly, for example, 60.76% flexibility area reduction for phase $\phi=\{b\}$; (ii) voltage unbalance limits have a major impact on flexibility services provided by units with a lack of coordination. It follows that a lack of phase coordination combined with voltage unbalance limits creates the worst conditions for providing flexibility services: single-phase units cannot effectively manage voltage unbalance across the network, which results in the infeasibility of services provision.
@@ III-B Case Study: 221-bus LV Distribution Network in the UK
To investigate the impact of phase unbalance on flexibility services, 12 single-phase flexible units (four units per phase) were placed in the network as shown in Fig. 7. Each unit has a P-Q capability of ±5 kW and ±5 kVAr. That is, units can independently regulate their active and reactive power output in the range of [-5.0,5.0] kW and [-5.0,5.0] kVAr. Then, the aggregated P-Q flexibility areas were calculated per each phase by iteratively solving the flexibility estimation model (1a)-(1p). The impact of phase unbalance and coordination on P-Q flexibility areas is shown in Fig. 8, where VUF constraints (1o) in the range 0.5-1.0% are imposed. [^5] To avoid scalability issues, in the 221-bus system, VUF constraints (1l)-(1o) are imposed only on the 7 buses with the highest voltage unbalances.
@@
Note that in this simulation setting, all flexible units have been assigned the same P-Q capabilities and distributed evenly across phases. Yet, their operation is limited considerably due to phase coordination issues and voltage unbalance. It follows that the provision of flexibility services by DER can be challenging even in LV networks with low load unbalance as flexible units can exacerbate existing voltage unbalances by producing or consuming power. To effectively manage such unbalances and voltage constraints, units connected to different phases must be perfectly coordinated. The simulations performed for the realistic 221-bus network demonstrate that lack of coordination between units significantly reduces the potential of DER flexibility services and can magnify voltage unbalance problems. Other aspects and implications of these findings are discussed in Section IV.
III-C Stochastic analysis
Previous simulations in this section focused on a deterministic analysis of aggregated P-Q flexibility in LV unbalanced networks. The deterministic setting enabled an interpretable analysis of the impacts of phase unbalance and DER coordination on flexibility provision, with clear visualisations. This subsection complements the reported findings by providing a stochastic analysis of voltage unbalances caused by unbalanced loads and flexible units. Specifically, the aim is to estimate how often VUF limit violations can occur when uncoordinated flexible resources operate in unbalanced networks. For this purpose, 1,000 load scenarios were randomly generated for the 221-bus LV distribution network. In each scenario, P-Q power demand values were varied independently across all buses by ±50% (sampled from a uniform distribution) to reflect temporal and spatial variability.
Then, to simulate the operation of flexible units, the following flexibility configurations were considered:
  • No flexibility provision: all units are switched off. This baseline solution corresponds to the intrinsic voltage imbalances in the network due to unbalanced loads.
  • Uncoordinated flexibility provision: all units operate independently and without coordination. Each unit can consume or generate power. The values of flexible power injections are generated randomly (uniformly) for each scenario within the capacity limits of the units.
  • Power consumption by phase A: only units connected to phase A are active, all consuming power. This stresses the system and exacerbates existing voltage unbalances in the network. The values of flexible loads are generated randomly (uniformly) for each scenario within the capacity limits of the units.
For each scenario, the OPF problem (1a)-(1k) was solved, and the maximum voltage unbalance across all buses was calculated. The results are presented in Fig. 9, where the VUF values are sorted in descending order, thus forming VUF duration curves across the scenarios. Note that without flexibility, unbalanced loads do not cause significant voltage unbalances – maximum VUF values are well below 1.3% for all scenarios. However, when introducing uncoordinated flexible resources, a significant share of scenarios (over 15%) exhibit violations of the voltage unbalance limit. These results highlight the importance of phase-aware control strategies for flexible resources: without coordination, flexibility provision can inadvertently introduce harmful voltage unbalances.
Figure 9: Sorted maximum VUF values (VUF duration curves) for 1,000 stochastic load scenarios under three flexibility configurations: (i) no flexibility provision, (ii) uncoordinated flexible power consumption/generation by all units, and (iii) power consumption only by units connected to phase A. The shaded red area highlights violations of the UK operational VUF limit of 1.3% for LV networks.
@@ IV Discussion
Multiple aspects of flexibility provision in unbalanced LV distribution networks have to be further investigated. One of the key factors affecting the aggregated P-Q flexibility is the relationship between the placement of flexible units and locations of voltage unbalances. For instance, in the 221-bus case study displayed in Fig. 7, the highest load unbalance is happening at the feeder with the flexible unit F5{a} connected. Six more flexible units are placed in this part of the network. Therefore, the design of this experiment inherently creates constraints on the operation of these flexible units due to the nearby voltage unbalances. That is, units must be coordinated when consuming or producing power to manage voltage constraints and keep the network’s operation feasible. For example, units connected to phase $\phi=\{a\}$ (the phase with the highest load) must not increase their power consumption while units in other phases consume power. Otherwise, voltage unbalance in some parts of the network may increase to unacceptable levels or even the voltage limits (1b) for phase $\{a\}$ may be violated. Note that this would not happen if flexible units were located far from the load unbalance, say, at the primary substation. Therefore, the question arises: how does the initial network unbalance impact the network’s P-Q flexibility and how does it restrict flexible resources at different locations? This question has implications for DER planning models. Depending on DSO metrics and objectives, some DER should be placed closer to the primary substation, while other flexible resources should be located closer to existing voltage unbalances to manage voltage constraints. A similar problem was considered in [23, 24], where it was demonstrated that suboptimal selection of the DER connection phase can lead to voltage unbalance and limit DER capacity.
@@ V Conclusion
This paper investigates the provision of flexibility services in unbalanced LV distribution networks and proposes a framework for quantifying the impacts of phase unbalance and imperfect coordination between distributed flexibility providers. At the core of the framework lies an accurate nonlinear three-phase AC OPF model designed for simulating the operation of flexible units connected to different phases. To characterise feasible flexibility services available in networks, the concept of aggregated P-Q flexibility areas is applied. The combination of these models and concepts allows to directly translate voltage unbalance and phase coordination constraints into reductions in aggregated P-Q flexibility. The simulations, performed for an illustrative 5-bus distribution system and a real 221-bus LV network in the UK, demonstrate that a significant share (over 30%) of total aggregated DER flexibility potential may be unavailable due to voltage unbalances and lack of coordination between DER connected to different phases.